April 22-25

– Four stores provided ~41.25 hours of luminosity
– Cryo-system personnel replaced a fly wheel and bearings on two Tevatron wet engines
– Linac personnel repaired a loose drive filament connection on LRF2
– D-Zero made several accesses for necessary repairs
– Linac personnel repaired a vacuum leak on RF tank #2
– Pbar worked on an Accumulator emittance problem
– The Tevatron quenched with no store in the machine

*The integrated luminosity for the period from 4/18/11 to 4/25/11 was 49.01 inverse picobarns. NuMI took 3.36E18 of beam during this same time period.

Read the Current Accelerator Update
Read the Early Bird Report
View the Tevatron Luminosity Charts